OpenStudy (anonymous):

Evaluate 64 × a-4 for a = 2. 4 8

OpenStudy (anonymous):

64(2-4)=

OpenStudy (anonymous):

2-4=-2 -2x64=-128

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

now what

OpenStudy (anonymous):

Now that's it. You said to evaluate it. It's been evaluated.

OpenStudy (anonymous):

Unless you are instructed to evaluate a=4 and a=8 as well

OpenStudy (anonymous):

im not sure whats the answer